Allied Electric Co, is developing a new design for its portable electric hair dryer Test market data indicate a demand for the new hair -dryer as follows:

Q-30,000-1000P

Where Q is hair dryer sales and P its price.

1. How many hair dryers could allied sell at Rs. 22.50 each? II. What price would Allied have to charge to sell 12,000 hair dryers?

III Calculate price elasticity of demand at price of Rs. 20
1
Expert's answer
2021-09-08T14:14:55-0400

(i)

Given

price 22.50

"Q=30,000-1000P"

Then

"Q=30000-1000(22.50)\\=30000-22500\\=7500"


(ii)

Given

Q=12000

"Q=30,000-1000P"

Then

"12000=30000-1000(P)\\P=3000012000\\P=18"


(ii)

Given

price 20

"Q=30,000-1000P"

Then

"Q=30000-1000(20)\\=30000-20000\\=10000"

price elasticity of demand"=\frac{\frac{Q_2-Q_1}{\frac{Q_2+Q_1}{2}}}{\frac{P_2-P_1}{\frac{P_2+P_1}{2}}}"

e(p)=price elasticity

Q1=old quantity demanded

Q2=new quantity demanded

P1=old price

P2=new price

"=\frac{\frac{10000-7500}{\frac{10000+7500}{2}}}{\frac{20-22.5}{\frac{20+22.5}{2}}}"

"=\frac{\frac{2500}{8750}}{\frac{2.5}{21.25}}"

"=\frac{0.286}{-0.118}\\=-2.42"